# ResearchTown: Simulator of human research community

### v1

@@ D.1 Data collection details
Author data collection. A primary challenge in collecting author data is that there might be multiple human researchers with the same name, and some human researchers may not have any publicly available publication records on public platforms including arXiv, Google Scholar, or Semantic Scholar. As a result, at the paper collection stage, we only have each author’s name. We use the semanticscholar Python package [^8] https://github.com/danielnsilva/semanticscholar to search for the author by name, verify that they have contributed to the specific target paper, and obtain a unique author ID from Semantic Scholar. This ID then allows us to retrieve their available publication information. To prevent information leakage when simulating paper writing and review scenarios, we exclude any of the author’s publications released after the target paper’s publication year. For example, if we target to simulate a paper published in 2022, we ignore all of the author’s publications appearing after 2022. We also exclude the target paper itself to avoid leaking information. Generally, we limit the maximum number of collected publications to around 20, focusing on those most relevant to the target time frame. Additionally, we gather each author’s co-author network and their top publications to enrich the dataset with useful relational information.
@@
D.2 PaperBench technical details
PaperBench is designed to evaluate the effectiveness of paper-writing simulations by gathering high-quality paper metadata from top-tier ML conferences, such as NeurIPS 2024 and ICLR 2024. Based on the collected author and paper data, we perform the following two post-processing steps:
First, we address cases where authors have no accessible publications beyond the current paper, or where citation data extraction fails due to API issues. In such cases, we exclude these papers. We only retain those with full author publication information, as well as complete metadata including introduction, abstract, title, and citations. After this filtering step, we end up with approximately 1,200 papers and then randomly sample 1,000 from them.
@@ D.2 PaperBench technical details
D.3 HighImpactPaperBench technical details
@@
For these impactful papers, it is crucial to prevent the inclusion of publications released after their publication year when gathering authors’ publication data. Such later works could significantly alter the researcher’s trajectory, misrepresent the historical context of these influential contributions, and leak information for simulation.
@@
After collecting paper and author data, we remove any papers with incomplete information due to crawling errors. From the remaining set, we randomly sample 100 papers to form the final benchmark. These selected papers have averaged over 100 citations in the past five years, ensuring that HighImpactPaperBench represents a collection of influential and well-established research.
D.4 Reviewbench technical details
After obtaining the reviewer, paper, and author data, we filter out any papers lacking valid reviews during crawling. From the remaining set, we randomly select 200 reviews each corresponding to one paper as ReviewBench.
Appendix E Ablation study technical details
@@
Data for paper-writing researcher number ablation.. Not all papers in ResearchBench have more than five authors. To ablate the effect of the number of researchers (1 to 5), we select a subset from the hard part of PaperBench within ResearchBench including 333 paper writing tasks, ensuring each paper has more than five authors. This filtering results in a subset of 172 paper-writing tasks. We focus on the hard subset because we believe that involving multiple research agents in more challenging scenarios yields a more significant difference in performance.
@@
Appendix F Evaluation technical details
In this section, we provide more implementation details related to our evaluation process.
Appendix G Additional experimental results
Table 3 includes detailed similarity score results for each question from Q1 to Q5, together with their average scores. Besides text-embedding-large-3 and voyage-3, we also include nv-embed-v2  [^11] https://huggingface.co/nvidia/NV-Embed-v2 as an additional state-of-the-art open-source embedding-based similarity metric.
Table 3: Detailed embedding-based similarity scores for PaperBench (Hard/Medium/Easy). We include three metrics (text-embedding-3-large, voyage-3, nv-embed-v2). Q1–Q5 are per-question similarity scores; “Avg” is their average.
Exp setting text-embedding-3-large voyage-3 nv-embed-v2
Q1 Q2 Q3 Q4 Q5 Avg Q1 Q2 Q3 Q4 Q5 Avg Q1 Q2 Q3 Q4 Q5 Avg
PaperBench-hard
Self-agg 31.69 48.89 48.64 43.51 42.69 43.08 65.11 53.89 49.67 48.09 47.15 52.78 38.03 46.67 41.81 44.05 40.75 42.26
Agent-agg 46.72 57.45 55.80 50.74 50.92 52.32 68.51 57.20 54.41 52.51 52.64 57.05 47.10 52.51 47.50 49.69 48.41 49.04
Data-agg 49.99 62.52 59.01 54.42 53.23 55.83 71.91 62.14 56.22 56.31 56.29 60.57 50.83 58.46 52.38 53.41 52.07 53.43
Global-agg 55.35 64.83 61.37 58.55 57.84 59.59 73.94 63.50 59.90 59.59 59.75 63.34 55.62 60.61 54.75 57.15 56.21 56.87
PaperBench-medium
Self-agg 32.77 49.60 48.86 43.78 43.00 43.60 64.96 54.09 49.08 47.96 46.88 52.60 38.66 47.37 42.04 43.92 41.17 42.63
Agent-agg 49.59 60.05 58.81 52.63 52.75 54.77 69.88 59.08 56.89 54.24 53.78 58.77 49.75 54.49 50.24 51.55 49.99 51.20
Data-agg 64.33 74.84 70.57 64.42 62.78 67.39 79.63 72.05 66.93 65.18 64.64 69.69 64.53 69.27 64.28 63.04 61.75 64.57
Global-agg 65.24 73.88 69.53 64.92 63.92 67.50 79.35 71.33 67.65 65.44 65.14 69.78 65.08 68.93 62.98 63.38 62.64 64.60
PaperBench-easy
Self-agg 33.78 50.05 48.95 44.65 43.90 44.26 65.59 54.49 49.14 48.70 47.93 53.17 39.82 47.81 42.40 44.72 42.48 43.44
Agent-agg 52.35 61.33 60.24 54.54 55.27 56.75 71.72 60.43 58.24 55.71 55.87 60.39 52.43 56.40 51.98 53.51 53.24 53.51
Data-agg 76.29 83.53 80.07 73.48 71.30 76.93 86.20 80.96 77.38 73.79 72.35 78.14 76.13 78.90 75.04 72.05 71.37 74.70
Global-agg 74.75 80.25 76.57 71.54 70.60 74.74 84.82 77.91 75.37 71.78 71.09 76.19 74.59 75.77 71.23 70.15 70.28 72.41
Appendix H Additional case study
@@
H.1 Additional case study for in-distribution evaluation
Tables 14, 15, 16, 17, and 18 present examples of tasks and their corresponding outputs for the in-distribution evaluation of ResearchTown. These examples illustrate the evaluation process defined in this work.
H.2 Additional case study for out-of-distribution application
In Table 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, we show examples of the inputs and outputs of the out-of-distribution application of ResearchTown. Additionally, each table caption includes a brief comment on the quality of the generated papers for reference.
@@ H.2 Additional case study for out-of-distribution application
System You are an autonomous intelligent agent tasked with writing the first-person persona of a research based on his publications.
@@
Publications - A list of paper abstracts written by the researcher that you will be writing of.
@@
You should focus more on recent publications, which reflect the researcher’s recent persona. You should be concise and clear. The persona should be ranging from 100 to 300 words.
@@
Learning node embeddings that capture a node’s position within the broader graph structure is crucial for many prediction tasks on graphs. However, existing Graph Neural Network (GNN) architectures have limited power in capturing the position/location of a given node concerning all other nodes of the graph. Here we propose Position-aware Graph Neural Networks (P-GNNs), a new class of GNNs for computing position-aware node embeddings. P-GNN first samples sets of anchor nodes computes the distance of a given target node to each anchor set, and then learns a non-linear distance-weighted aggregation scheme over the anchor sets. This way P-GNNs can capture the positions/locations of nodes concerning the anchor nodes. P-GNNs have several advantages: they are inductive, scalable, and can incorporate node feature information. We apply P-GNNs to multiple prediction tasks including link prediction and community detection. We show that P-GNNs consistently outperform state-of-the-art GNNs, with up to 66% improvement in terms of the ROC AUC score.
@@
AutoML has demonstrated remarkable success in finding an effective neural architecture for a given machine learning task defined by a specific dataset and an evaluation metric. However, most present AutoML techniques consider each task independently from scratch, which requires exploring many architectures, leading to high computational costs. Here we propose AutoTransfer, an AutoML solution that improves search efficiency by transferring the prior architectural design knowledge to the novel task of interest. Our key innovation includes a task-model bank that captures the model performance over a diverse set of GNN architectures and tasks and a computationally efficient task embedding that can accurately measure the similarity among different tasks. Based on the task-model bank and the task embeddings, we estimate the design priors of desirable models of the novel task, by aggregating a similarity-weighted sum of the top-K design distributions on tasks that are similar to the task of interest. The computed design priors can be used with any AutoML search algorithm. We evaluate AutoTransfer on six datasets in the graph machine learning domain. Experiments demonstrate that (i) our proposed task embedding can be computed efficiently and that tasks with similar embeddings have similar best-performing architectures; (ii) AutoTransfer significantly improves search efficiency with the transferred design priors, reducing the number of explored architectures by an order of magnitude. Finally, we release GNN-Bank-101, a large-scale dataset of detailed GNN training information of 120,000 task-model combinations to facilitate and inspire future research.
@@
Here is a high-level summarized insight of the Machine Learning research field.
@@
User Here is a high-level summarized insight of a research field: Machine Learning.
@@
Strength - Advantages and strengths of the submission that can improve its chances to be accepted.
@@
Clarity: Is the writing clear, structured, and terms defined?
@@
Score - A score between 1 to 10 to evaluate the overall quality of the submission to an academic journal. It should be one of 1, 2, …, 10. 1 is the lowest score while 10 is the highest score.
@@
Please evaluate the submission based on the summarized strengths and weaknesses provided. The score should be more related to weakness. If there is a critical weakness in the submission, you should give a lower score. If the submission has a minor weakness, you can give a higher score. If the submission has no weakness, you should give a high score. But the strengths should also be considered in the evaluation.
@@
8/10: The submission is among the top 10% of all the papers. It provides sufficient justification for all its arguments and claims. Some extra experimentation is needed, but they are not essential. The proposed method is very original and can generalize to various fields. This submission deepens the understanding of some phenomena, or lowers the bar for future research on an existing problem.
@@
Reviews - It typically contains the score, weakness, and weakness of the submission, each by a different reviewer.
@@
User Here is a high-level summarized insight of a research field of machine learning.
@@
Discuss how such paper will affect the future research.
@@
Outline your proposed methodology in detail, including the method, dataset, metric that you plan to use. Describe the expected outcomes.
@@
Please provide the five core questions contents based on the above content.
@@
Why is it hard? The challenges in solving this problem stem from the inherent complexity of adversarial attacks, which can manipulate model predictions in subtle ways. Naive approaches may fail because they do not account for the diverse nature of adversarial perturbations, leading to inadequate coverage guarantees. Additionally, the computational overhead associated with randomized smoothing techniques complicates the implementation of robust conformal prediction, as it requires extensive sampling and can significantly increase training time. Overcoming these technical and practical obstacles is essential to develop an effective solution. The challenge arises from the complexities associated with label noise and adversarial perturbations, which can distort data distributions and violate the assumptions of traditional conformal prediction methods. Existing approaches often fail to account for the adversarial nature of noise or the distribution shifts that occur during inference. Additionally, ensuring valid coverage guarantees while maintaining model performance requires sophisticated techniques that balance robustness and accuracy, complicating the design of effective algorithms.
@@
- Chapter 3 performs a manual analysis of 500 sample images from the LAION dataset to categorize them based on the correlation between image features and captions but lacks metrics to quantify the representativeness of this sample within the entire dataset.

### v2

@@ E.1 Data collection details
Author data collection. A primary challenge in collecting author data is that there might be multiple human researchers with the same name, and some human researchers may not have any publicly available publication records on public platforms, including arXiv, Google Scholar, or Semantic Scholar. As a result, at the paper collection stage, we only have each author’s name. We use the semanticscholar Python package [^10] https://github.com/danielnsilva/semanticscholar to search for the author by name, verify that they have contributed to the specific target paper, and obtain a unique author ID from Semantic Scholar. This ID then allows us to retrieve their available publication information. To prevent information leakage when simulating paper writing and review scenarios, we exclude any of the author’s publications released after the target paper’s publication year. For example, if we aim to simulate a paper published in 2022, we ignore all of the author’s publications appearing after 2022. We also exclude the target paper itself to avoid leaking information. Generally, we limit the maximum number of collected publications to around 20, focusing on those most relevant to the target time frame. Additionally, we gather each author’s co-author network and their top publications to enrich the dataset with useful relational information.
@@
E.2 PaperBench details
PaperBench is designed to evaluate the effectiveness of paper-writing simulations by gathering high-quality paper metadata from top-tier ML conferences, such as NeurIPS 2024 and ICLR 2024. Both NeurIPS 2024 and ICLR 2024 post-date beyond GPT-4o-mini’s October 2023 knowledge cutoff. Thus, data leakage is not a concern. We also mask the full text during the simulation to avoid accidental exposure. Based on the collected author and paper data, we perform the following two post-processing steps:
First, we address cases where authors have no accessible publications beyond the current paper or where citation data extraction fails due to API issues. In such cases, we exclude these papers. We only retain those with full author publication information, as well as complete metadata including introduction, abstract, title, and citations. After this filtering step, we end up with approximately 1,200 papers, and then randomly sample 1,000 from them.
@@ E.2 PaperBench details
E.3 Reviewbench details
After obtaining the reviewer, paper, and author data, we filter out any papers lacking valid reviews during crawling. From the remaining set, we randomly select 200 reviews, each corresponding to one paper as ReviewBench.
E.4 HighImpactPaperBench details
@@
For these impactful papers, it is crucial to prevent the inclusion of publications released after their publication year when gathering authors’ publication data. Later works such as these could significantly alter the trajectory of the researcher, misrepresent the historical context of these influential contributions, and leak information for simulation. After collecting paper and author data, we remove any papers with incomplete information due to crawling errors. From the remaining set, we randomly sample 100 papers to form the final benchmark. These selected papers have averaged over 100 citations in the past five years, ensuring that HighImpactPaperBench represents a collection of influential and well-established research.
The motivation for using impactful papers as evaluation is to use them as an extreme-case test for idea simulation. While some may exist in the LLM’s training data, this benchmark is separate from our main results and serves to explore how LLMs handle well-known concepts. Our similarity analysis shows that 55% of generated papers score between 0.65–0.75, and 18% exceed 0.75, indicating moderate to high alignment. Only 1% scored below 0.45. These scores are comparable to PaperBench, suggesting no abnormal inflation. Even famous papers like VAE, GAN, and LayerNorm do not receive notably high scores, implying that semantic similarity, not memorization based on citation relationships, drives the results, especially for tool/benchmark papers, which naturally resemble their references more.
Appendix F Embedding-based Evaluation Details
In this section, we first explain the motivation for our designed multi-component embedding-based evaluation, then we provide a more formal definition and implementation details related to our evaluation process.
Decompositionality. A single idea or a review can manifest through diverse descriptions or implementation strategies. Therefore, directly applying a cosine similarity-based metric is inadequate for capturing conceptual equivalence. To solve this, we design point-wise evaluation metrics to paraphrase the paper and review it into aligned key points with the same LLM-based prompting. This structure enables alignment between papers that differ methodologically but share similar motivations and problem framings. For instance, in Chen et al. (2025) and Jin et al. (2025), despite distinct methods and settings, experts would find strong alignment on the motivations and core concepts in these papers.
Scalability. To address the challenge that a single idea can take many concrete forms, we complement decomposition with scalability. LLMs can generate hundreds of semantically distinct research questions from a single prompt, but evaluating these outputs traditionally requires domain experts—a process that is costly, slow, unscalable, and hard to reproduce. For example, Si et al. (2024) spent thousands hiring top-tier researchers solely for annotation and review, which is infeasible for evaluating large-scale, automated research generation. Our approach replaces this bottleneck with semantic similarity over 5Q-decomposed representations. We can select the best among the samples and make the score the final result.
Extensibility. While we acknowledge the importance of elements like mathematical formulations or algorithmic workflows, our framework is inherently extensible—the original format can be expanded with more key points by adding domain-specific dimensions such as algorithmic structure or key theoretical results. This is especially valuable in systems and theory papers, enabling more fine-grained and domain-aware similarity analysis. As demonstrated in [Fine-Grained Evaluation with LLM and Human], our approach also supports the integration of non-semantic metrics like logical consistency and factual accuracy, making it extensible from an evaluation metric perspective.
Reliability. Our embedding-based / LLM-based similarity metric builds on state-of-the-art models optimized for knowledge-intensive tasks. Voyage AI embeddings, widely adopted in real-world RAG systems, are designed to reduce hallucination and excel in high-precision semantic retrieval, making them ideal for evaluating research content. Additionally, state-of-the-art LLMs are highly effective at semantic comparison.
Baselines for evaluation. To check whether ResearchTown provides a realistic simulation, we benchmark similarity in real-world research activity. For paper writing, we reference two concurrent papers (Chen et al., 2025; Jin et al., 2025) recognized for presenting nearly identical ideas, yet with different writing styles and experiments, which yield a VoyageAI similarity of 0.8244. This suggests that scores above 0.82 can potentially indicate strong idea overlap. For review writing, we analyze the data of reviewers evaluating the same paper. The average inter-reviewer similarity is 0.5900 (strengths) and 0.5904 (weaknesses), reflecting natural variance in human judgment. These inter-similarity scores in the real world confirm that similarity scores represent realistic simulation.
where $\textsc{sim}(\cdot,\cdot)$ represents an embedding-based similarity metric, such as voyage-3  [^14] https://blog.voyageai.com/2024/09/18/voyage-3/ and text-embedding-large-3  [^15] https://openai.com/index/new-embedding-models-and-api-updates/. By leveraging the LLM to generate structured embeddings for each question, this approach ensures a meaningful and consistent comparison of the generated and ground-truth papers.
where $\textsc{sim}(\cdot,\cdot)$ refers to the same similarity metric in paper writing evaluation. This metric emphasizes the recall rate of the generated review by measuring whether each point in the real-world review is potentially included in the generated review. Since each review consists of both strengths and weaknesses, we compute separate similarity scores for strengths and weaknesses. Additionally, since both $\mathbf{r}_{v}$ and $\mathbf{r}_{v}^{*}$ include a final score $\mathbf{S}_{v}$ and $\mathbf{S}^{*}_{v}$ as attributes, we calculate $\Delta\mathbf{S}_{v}=|\mathbf{S}_{v}-\mathbf{S}_{v}^{*}|$ to quantify the difference between the generated and real-world review scores.
Appendix G LLM-based Evaluation Details
In this section, we provide more technical details about using LLM prompting for evaluation.
Prompting for similarity. For prompting-based evaluation, we decompose overall similarity into six fine-grained dimensions: (1) topic consistency, (2) method consistency, (3) factual consistency, (4) claim consistency, (5) application context consistency, and (6) overall semantic similarity. These dimensions are designed to capture distinct yet complementary aspects of alignment between the generated and reference proposals, ranging from high-level research focus (such as topic and application context) to specific technical content (such as methods, facts, and claims). Importantly, they are intended to capture nuances that may not be easily detected by embedding-based models, enabling a more comprehensive and interpretable assessment than relying on a single similarity score. Each dimension is rated on a scale from 0 to 10.
Prompting for novelty and feasibility. In addition to measuring similarity, we prompt LLMs to assess two intrinsic quality dimensions: (1) novelty and (2) feasibility, which we consider essential characteristics of a strong research proposal. While similarity captures how well the generated content aligns with a reference, it does not fully reflect the proposal’s originality or practicality. These intrinsic dimensions address that gap by evaluating the creativity of the proposed idea and its potential for real-world implementation. Each dimension is scored on a scale from 0 to 10, complementing similarity-based metrics for a more holistic evaluation.
Prompt. To enable efficient evaluation, we adopt parallel prompting, where both the reference and generated proposals are input to the LLM in a single prompt, along with all evaluation criteria. This allows the model to produce scores for all dimensions simultaneously in one forward pass. The detailed descriptions of these evaluation criteria and the full prompts are provided in Table 17.
Appendix H Human Evaluation Details
Annotator Information. We recruit two graduate-level students with backgrounds in computer science and artificial intelligence. Both annotators have prior experience publishing in top-tier machine-learning conferences.
Annotated Data. We randomly sample 40 reference proposals and their corresponding generated proposals from PaperBench for human evaluation. We ask annotators to annotate on overall similarity, novelty, and feasibility.
Annotation Process. The annotation process consists of three stages: (1) preliminary annotation, (2) discussion, and (3) final annotation. In the preliminary stage, each annotator independently labels 10 examples. They then meet to discuss discrepancies, align their understanding, and refine the annotation criteria. Based on this discussion, they proceed to annotate the official 40 examples using the agreed-upon guidelines as the final results.
Annotation Instructions. At the start, annotators receive the same input information as used in the LLM-based prompting setup. During the discussion phase, they collaboratively develop more detailed and consistent annotation guidelines to ensure alignment in their final evaluations.
Appendix I Ablation Study Details
@@
Data for paper-writing researcher number ablation.. Not all papers in ResearchBench have more than five authors. To ablate the effect of the number of researchers (1 to 5), we select a subset from the hard part of PaperBench within ResearchBench, including 333 paper writing tasks, ensuring each paper has more than five authors. This filtering results in a subset of 172 paper-writing tasks. We focus on the hard subset because we believe that involving multiple research agents in more challenging scenarios yields a more significant difference in performance.
@@
Appendix J Additional Experimental Results
We provide more comprehensive experimental results on each sub-part of ResearchBench (PaperBench, ReviewBench, HighImpactPaperBench) in this section.
Additional Results on PaperBench. Table 5 shows that all models—Qwen-2.5-7B-Instruct, GPT-4o-mini, and Deepseek-v3—consistently achieve better performance with richer reference contexts (AGG-data and AGG-global) compared to narrower ones (AGG-self and AGG-agent), highlighting the importance of contextual information in similarity evaluation.
Additional Results on ReviewBench. As shown in Table 6, voyage-3 embeddings yield higher strength scores and larger $\Delta\mathbf{S}$ values than text-embedding-3, indicating greater discriminative power. While Qwen-2.5-7B-Instruct maintains strong similarity scores across all aggregation types, it exhibits larger deviations from human scores, suggesting potential scoring bias or overconfidence in its own outputs.
Additional results on HighImpactPaperBench. Besides the full results on PaperBench, we also evaluate ResearchTown under extreme conditions by attempting to simulate 100 of the most-cited machine learning papers from the past decade. ResearchTown achieves low similarity scores for papers introducing groundbreaking methods, such as “Layer Normalization” (Ba, 2016), or novel topics, such as “Energy and Policy Considerations for Deep Learning in NLP” (Strubell et al., 2019). However, the framework performs notably better on impactful papers focused on analysis or tool development. For instance, it achieves a similarity score exceeding 0.8 for papers like “Is BERT Really Robust? A Strong Baseline for Natural Language Attack on Text Classification and Entailment” (Jin et al., 2020), which provides adversarial analysis, and “Stanza: A Python Natural Language Processing Toolkit for Many Human Languages” (Qi et al., 2020), which offers a practical toolkit. These results suggest that high-impact research ideas may be more feasible than commonly perceived, and ResearchTown could potentially serve as a tool to inspire future impactful research.
Table 5: Evaluation results on embedding-based similarity score of PaperBench. We include comprehensive results on three different models and include sub-scores from Q1-Q5.
Model Component AGG-self AGG-agent AGG-data AGG-global
easy mid hard easy mid hard easy mid hard easy mid hard
DeepSeek-v3 Q1 43.84 44.08 41.74 52.46 49.93 47.33 68.18 61.70 54.74 66.10 61.20 55.66
Q2 51.52 51.71 50.02 61.74 59.75 58.29 75.53 69.68 63.92 74.29 69.81 65.76
Q3 50.84 51.44 49.58 59.74 58.57 56.32 71.02 67.05 61.03 70.32 67.18 62.33
Q4 48.51 48.87 47.80 56.21 54.09 53.22 67.55 63.11 59.86 66.31 62.72 60.01
Q5 50.29 50.02 49.03 60.37 58.44 56.43 68.04 64.38 59.96 69.51 65.82 62.90
Overall 49.00 49.22 47.64 58.10 56.16 54.32 70.06 65.18 59.90 69.31 65.35 61.33
gpt-4o-mini Q1 36.93 35.88 35.80 53.51 51.70 48.62 70.98 61.79 49.26 72.78 65.22 56.62
Q2 52.82 52.44 52.22 61.74 60.30 58.24 81.14 73.34 62.54 79.16 73.55 66.20
Q3 50.10 50.16 49.85 59.09 58.65 55.69 76.58 69.12 58.13 74.40 69.20 61.47
Q4 45.71 44.82 44.87 54.97 53.23 51.73 71.37 63.03 54.33 71.13 64.92 59.22
Q5 46.54 46.32 46.76 55.18 53.85 52.03 71.71 64.84 55.84 71.46 66.35 60.93
Overall 46.42 45.92 45.90 56.90 55.55 53.26 74.36 66.42 56.02 73.79 67.85 60.89
Qwen-2.5-7B-Instruct Q1 40.41 39.90 40.05 49.58 48.91 45.11 70.34 63.38 57.73 71.51 64.55 58.52
Q2 52.67 52.43 53.00 61.74 61.40 58.10 75.17 71.08 66.74 75.16 71.22 66.81
Q3 50.34 50.39 50.66 58.55 57.82 54.56 73.60 67.96 63.50 73.40 68.09 63.62
Q4 42.58 42.24 42.65 50.97 49.99 47.35 63.66 58.18 54.98 62.56 57.62 53.80
Q5 46.46 46.11 46.86 56.66 55.26 52.64 66.78 63.35 58.97 68.13 64.54 60.07
Overall 46.49 46.21 46.65 55.50 54.68 51.55 69.91 64.79 60.38 70.15 65.20 60.56
Table 6: Evaluation results on embedding-based similarity score of ReviewBench. We include comprehensive results on three different models and include strengths, weaknesses, and $\Delta\mathbf{S}$.
Setting Model text-embedding-3 voyage-3
Weakness Strength $\Delta S$ Weakness Strength $\Delta S$
AGG-self Qwen-2.5-7B-Instruct 49.79 52.08 1.36 65.24 64.82 1.36
Deepseek-v3 48.92 51.98 1.11 62.41 65.18 1.11
gpt-4o-mini 47.16 51.23 1.27 61.24 65.18 1.27
AGG-agent Qwen-2.5-7B-Instruct 50.12 51.38 1.41 66.24 65.49 1.41
Deepseek-v3 48.56 51.68 1.05 62.80 65.38 1.05
gpt-4o-mini 46.75 51.66 1.19 61.29 66.03 1.19
AGG-data Qwen-2.5-7B-Instruct 50.26 51.66 1.28 66.09 65.05 1.28
Deepseek-v3 49.21 51.34 1.07 63.11 65.19 1.07
gpt-4o-mini 47.62 51.45 1.26 61.74 65.57 1.26
AGG-global (k=5) Qwen-2.5-7B-Instruct 50.21 50.78 0.79 65.58 63.72 0.79
Deepseek-v3 48.95 50.57 0.81 62.56 64.21 0.81
gpt-4o-mini 51.51 47.17 1.55 66.01 61.39 1.55
Appendix K Additional Case Study
@@
Additional case study for in-distribution evaluation. Tables 18, 19, 20, 21, and 22 present examples of tasks and their corresponding outputs for the in-distribution evaluation of ResearchTown. These examples illustrate the evaluation process defined in this work.
Additional case study for out-of-distribution application. In Table 23, 24, 25, 26, 27, 28, 29, 30, 31, and 32, we show examples of the inputs and outputs of the out-of-distribution application of ResearchTown. Additionally, each table caption includes a brief comment on the quality of the generated papers for reference.
@@
System You are an autonomous intelligent agent tasked with writing the first-person persona of a researcher based on his publications.
@@
Publications - A list of paper abstracts written by the researcher that you will be writing about.
@@
You should focus more on recent publications, which reflect the researcher’s recent persona. You should be concise and clear. The persona should range from 100 to 300 words.
@@
Learning node embeddings that capture a node’s position within the broader graph structure is crucial for many prediction tasks on graphs. However, existing Graph Neural Network (GNN) architectures have limited power in capturing the position/location of a given node concerning all other nodes of the graph. Here we propose Position-aware Graph Neural Networks (P-GNNs), a new class of GNNs for computing position-aware node embeddings. P-GNN first samples sets of anchor nodes, computes the distance of a given target node to each anchor set, and then learns a non-linear distance-weighted aggregation scheme over the anchor sets. This way, P-GNNs can capture the positions/locations of nodes concerning the anchor nodes. P-GNNs have several advantages: they are inductive, scalable, and can incorporate node feature information. We apply P-GNNs to multiple prediction tasks, including link prediction and community detection. We show that P-GNNs consistently outperform state-of-the-art GNNs, with up to 66% improvement in terms of the ROC AUC score.
@@
AutoML has demonstrated remarkable success in finding an effective neural architecture for a given machine learning task defined by a specific dataset and an evaluation metric. However, most present AutoML techniques consider each task independently from scratch, which requires exploring many architectures, leading to high computational costs. Here we propose AutoTransfer, an AutoML solution that improves search efficiency by transferring the prior architectural design knowledge to the novel task of interest. Our key innovation includes a task-model bank that captures the model performance over a diverse set of GNN architectures and tasks, and a computationally efficient task embedding that can accurately measure the similarity among different tasks. Based on the task-model bank and the task embeddings, we estimate the design priors of desirable models of the novel task by aggregating a similarity-weighted sum of the top-K design distributions on tasks that are similar to the task of interest. The computed design priors can be used with any AutoML search algorithm. We evaluate AutoTransfer on six datasets in the graph machine learning domain. Experiments demonstrate that (i) our proposed task embedding can be computed efficiently and that tasks with similar embeddings have similar best-performing architectures; (ii) AutoTransfer significantly improves search efficiency with the transferred design priors, reducing the number of explored architectures by an order of magnitude. Finally, we release GNN-Bank-101, a large-scale dataset of detailed GNN training information of 120,000 task-model combinations to facilitate and inspire future research.
@@
Here is a high-level, summarized insight into the Machine Learning research field.
@@
User Here is a high-level, summarized insight of a research field: Machine Learning.
@@
Strength - Advantages and strengths of the submission that can improve its chances of being accepted.
@@
Clarity: Is the writing clear, structured, and do terms defined?
@@
Score - A score between 1 to 10 to evaluate the overall quality of the submission to an academic journal. It should be one of 1, 2, …, 10. 1 is the lowest score, while 10 is the highest score.
@@
Please evaluate the submission based on the summarized strengths and weaknesses provided. The score should be more related to weakness. If there is a critical weakness in the submission, you should give a lower score. If the submission has a minor weakness, you can give a higher score. If the submission has no weaknesses, you should give a high score. But the strengths should also be considered in the evaluation.
@@
8/10: The submission is among the top 10% of all the papers. It provides sufficient justification for all its arguments and claims. Some extra experimentation is needed, but they are not essential. The proposed method is very original and can be generalized to various fields. This submission deepens the understanding of some phenomena or lowers the bar for future research on an existing problem.
@@
Reviews - It typically contains the score, weakness, and strength of the submission, each by a different reviewer.
@@
User Here is a high-level, summarized insight into a research field of machine learning.
@@
Discuss how such a paper will affect future research.
@@
Outline your proposed methodology in detail, including the method, dataset, and metric that you plan to use. Describe the expected outcomes.
@@
Please provide the five core questions based on the above content.
@@
Role Content
User You are an expert evaluator of research proposals. Given a reference and a generated proposal, rate their similarity from 1 (completely different) to 10 (nearly identical) for each of the following dimensions:
- Topic Consistency: Do both proposals address the same research topic or problem area?
- Method Consistency: Are the research methods and approaches used in both proposals similar?
- Factual Consistency: Are the datasets, metrics, and models mentioned consistent between the two proposals?
- Claim Consistency: Do both proposals present similar conclusions or findings?
- Application Context Consistency: Are the application domains or use-cases targeted by both proposals the same?
- Overall Semantic Similarity: Do the two proposals convey the same overall idea or message?
Additionally, assess the following for each proposal individually:
- Novelty of Reference Proposal: Does the reference proposal introduce new ideas or perspectives?
- Feasibility of Reference Proposal: How realistic and implementable is the reference proposal based on its described methods and objectives?
- Novelty of Generated Proposal: Does the generated proposal introduce new ideas or perspectives?
- Feasibility of Generated Proposal: How realistic and implementable is the generated proposal based on its described methods and objectives?
Respond strictly in JSON format like:
  ”Topic Consistency”: X,
  ”Method Consistency”: X,
  ”Factual Consistency”: X,
  ”Claim Consistency”: X,
  ”Application Context Consistency”: X,
  ”Overall Semantic Similarity”: X,
  ”Novelty of Reference Proposal”: X,
  ”Feasibility of Reference Proposal”: X,
  ”Novelty of Generated Proposal”: X,
  ”Feasibility of Generated Proposal”: X
Table 17: Prompt for LLM-based evaluation of fine-grained similarity and quality assessment of research proposals.
@@
Why is it hard? The challenges in solving this problem stem from the inherent complexity of adversarial attacks, which can manipulate model predictions in subtle ways. Naive approaches may fail because they do not account for the diverse nature of adversarial perturbations, leading to inadequate coverage guarantees. Additionally, the computational overhead associated with randomized smoothing techniques complicates the implementation of robust conformal prediction, as it requires extensive sampling and can significantly increase training time. Overcoming these technical and practical obstacles is essential to developing an effective solution. The challenge arises from the complexities associated with label noise and adversarial perturbations, which can distort data distributions and violate the assumptions of traditional conformal prediction methods. Existing approaches often fail to account for the adversarial nature of noise or the distribution shifts that occur during inference. Additionally, ensuring valid coverage guarantees while maintaining model performance requires sophisticated techniques that balance robustness and accuracy, complicating the design of effective algorithms.
@@
- Chapter 3 performs a manual analysis of 500 sample images from the LAION dataset to categorize them based on the correlation between image features and captions, but lacks metrics to quantify the representativeness of this sample within the entire dataset.